# FlexVAR: Flexible Visual Autoregressive Modeling without Residual Prediction

### v1

@@ Abstract
This work challenges the residual prediction paradigm in visual autoregressive modeling and presents FlexVAR, a new Flexible Visual AutoRegressive image generation paradigm. FlexVAR facilitates autoregressive learning with ground-truth prediction, enabling each step to independently produce plausible images. This simple, intuitive approach swiftly learns visual distributions and makes the generation process more flexible and adaptable. Trained solely on low-resolution images ($\leq$ 256px), FlexVAR can: (1) Generate images of various resolutions and aspect ratios, even exceeding the resolution of the training images. (2) Support various image-to-image tasks, including image refinement, in/out-painting, and image expansion. (3) Adapt to various autoregressive steps, allowing for faster inference with fewer steps or enhancing image quality with more steps. Our 1.0B model outperforms its VAR counterpart on the ImageNet 256$\times$256 benchmark. Moreover, when zero-shot transfer the image generation process with 13 steps, the performance further improves to 2.08 FID, outperforming state-of-the-art autoregressive models AiM/VAR by 0.25/0.28 FID and popular diffusion models LDM/DiT by 1.52/0.19 FID, respectively. When transferring our 1.0B model to the ImageNet 512$\times$512 benchmark in a zero-shot manner, FlexVAR achieves competitive results compared to the VAR 2.3B model, which is a fully supervised model trained at 512$\times$512 resolution.
@@
Figure 1: Generated samples from FlexVAR-$d24$ (1.0B). FlexVAR generates images with various resolutions and aspect ratios, even though it is trained with a resolution of $\leq$ 256$\times$256.
@@ 1 Introduction
Autoregressive (AR) models aim to learn the probability distribution of the next token, offering great flexibility by generating tokens of any length. This design brings significant advancements in the field of Natural Language Processing (NLP), demonstrating satisfactory generality and transferability [4, 30, 31]. Concurrently, the computer vision field has been striving to develop large autoregressive models [27, 28, 1, 43, 29, 38]. These models employ visual tokenizers to discretize images into a series of 1D tokens [34, 22, 49, 52, 41] or 2D scales [44, 51, 42, 37, 25] and then utilize AR to model the next unit. However, these image autoregressive models typically output images at a single resolution, the flexibility of AR has not yet been realized.
Recently, in image generation, VAR [44] has pioneered scale-wise autoregressive modeling, completing image autoregression based on 2D sequences. This approach predicts the next scale rather than the next token, thereby preserving the 2D structure of images and mitigating the issue of limited receptive fields in 1D causal transformers. Specifically, VAR predicts the ground-truth (GT) [^1] To avoid confusion, we use ground-truth (GT) to represent image latent feature, and residual to represent residual latent feature. of the smallest scale in the first step. Subsequently, at each step, it predicts the residuals of the current scale and the prior one. Finally, the outputs of each scale are upsampled to a uniform size and undergo weighted summation to generate the final output, as illustrated in Fig. 2(a). Successors [51, 42, 37, 25] have all adopted the residual design, assuming it to be effective. Although this technique achieves commendable performance, it encounters a primary challenge: The residual prediction relies on a rigid step design, restricting the flexibility to generate images with varying resolutions and aspect ratios, thus limiting the adaptability and flexibility of image generation. Meanwhile, residuals at different scales often lack semantic continuity, and this implicit prediction approach may limit the model’s capacity to represent diverse image variations.
@@
Figure 2: Comparison between VAR [44] and our FlexVAR. VAR predicts the GT1 in step 1 and then predicts the residuals relative to the GT in all subsequent steps. Our FlexVAR predicts the GT at each step.
@@ 2 Related Work
VQ-VAE [46] introduces a groundbreaking two-stage image generation paradigm: (1) encoding the image into a latent space and quantizing it to the nearest code in a fixed-size codebook; (2) modeling the discretized code using PixelCNN [45], which predicts the probability distribution of each code in raster scan order. This two-stage paradigm has laid the foundation for many subsequent works.
@@
Raster-scan Manner Building on the aforementioned foundation, [12, 33] perform autoregressive learning in latent space with Transformer architecture. VQVAE-2 [35] and RQ-Transformer [22] use extra scales or stacked codes for next-image-token prediction. These works further advance the field and achieve impressive results. Recently, [41, 26] utilize a GPT-style next-token-prediction strategy to achieve high-quality image generation. [17] further improves this paradigm by introducing a mixture of autoregressive models, while [23] incorporates Mamba structure [13] to accelerate image generation. [48, 53, 14] combine diffusion processes into autoregressive modeling to address the information loss caused by quantization, which potentially degrades the quality of generated images.
@@
Random-scan Manner Masked-prediction models learn to predict masked tokens in a BERT-style manner [9, 16, 2]. They introduce a bidirectional transformer that predicts masked tokens by attending to unmasked conditions, thus generating image tokens in a random-scan manner. This approach enables parallel token generation at each step, significantly improving inference efficiency. Specifically, [5, 6] apply masked-prediction models in class-to-image and text-to-image generation, respectively. MagViT series [50, 29] adapts this approach to videos by introducing a VQVAE for both images and videos. NOVA [8] first predicts temporal frames and then predicts spatial sets within each frame to achieve high-quality image/video generation.
@@
Scaling-scan Manner VAR [44] establishes a new generation paradigm that redefines autoregressive learning on images from next-token-prediction to next-scale-prediction. VAR in parallel predicts image tokens at one scale, significantly reducing the number of inference steps. Following VAR, VAR-CLIP [51] achieves text-to-image generation by converting the class condition token into text tokens obtained from the CLIP Text Encoder. In terms of operational efficiency, [7] introduces an efficient decoding strategy, [37] incorporates linear attention mechanisms to accelerate image generation, and [25] designs a lightweight image quantizer, significantly reducing training costs. Regarding generation quality, [42, 36] optimize image details by using continuous tokenizers in combination with flow matching or diffusion model. Infinity [15] redefines the visual autoregressive model under a bitwise token prediction framework, remarkably enhancing generation capability and detail.
@@ 3.1 Scale-wise Autoregression (Preliminary)
Previous approaches [44, 51, 7, 37, 25, 42] typically follow a residual prediction paradigm. They only regress the ground-truth at the first scale ($g_{1}$), while at subsequent $i^{th}$ scale, the residual between the preceding scale ($g_{i-1}$) and the current scale ($g_{i}$) is predicted. We formulate these residuals ($\{r_{i}\}^{n}_{i=2}$) as :
@@
here $\mathrm{Upsample^{i}}$ represents upsample $g_{i-1}$ to the $i^{th}$ scale. The autoregressive likelihood is formulated as:
@@
attention mechanisms (e.g., Transformer [47]) are utilized to instantiate this modeling. During the $i^{th}$ autoregressive step, all preceding residuals are merged in the autoregressive model, which then predicts the probability distribution of $r_{i}$. The $h_{i}\times w_{i}$ tokens in $r_{i}$ are generated in parallel, conditioned on all preceding units. Thus, image token maps can be redefined as: $T=\{g_{1},r_{2},r_{3}...,r_{n}\}$. Finally, each image token map in $T$ is upsampled to $\mathcal{R}^{h_{n}\times w_{n}}$ and summarized for image generation.
@@ 3.3 Quantize & reconstruct images at various scales
Encoding. Given an input image $I\in\mathcal{R}^{H\times W}$, an autoencoder $\mathcal{E}(\cdot)$ [12] is used to convert $I$ into latent space $f$:
@@
here $h=\frac{H}{16},w=\frac{W}{16}$. We then downsample $f$ at $K$ random scales to obtain multi-scale latent features =$\mathcal{F}=\{f_{1},f_{2},...,f_{K}\}$. $f_{k}$ represents represents the latent feature of the $k^{th}$ downsample from $f$. $f_{K}$ matches the original resolution of $f$.
@@
Decoding. The multi-scale images ^=$\hat{\mathcal{I}}=\{\hat{I_{1}},\hat{I_{2}},...,\hat{I_{K}}\}$ are reconstructed using the decoder $\mathcal{D}(\cdot)$ [12] given $q^{(k,i,j)}$. We follow Llamagen [41] to adopt the same loss functions ($\mathcal{L}_{vae}$) to train $\{\mathcal{E},\mathcal{Q},\mathcal{D}\}$ at each scale without special design. Therefore, the final loss function can be formulated:
@@ 3.4 Visual autoregressive modeling without residual
In our FlexVAR, we design a 2D scalable PE ($\mathcal{P}\in\mathcal{R}^{d\times 2h\times 2w}$) adding to the h-w coordinates. It contains $2h\times 2w$ learnable queries with $d$ channels. At the ith step, $\mathcal{P}$ is upsampled/downsampled to match the scale of $g_{i}$. To ensure stability during linear interpolation across various scales, we set $\mathcal{P}$ to 2×2$\times$2 × the size of the max latent space in traning. $\mathcal{P}$ is initialized using 2D sin-cosine PE [11] to ensure the 2D positional correlation. Additionally, we experimentally find that in our ground-truth prediction paradigm, incorporating PE for step embeddings is unnecessary (Tab. 6). Therefore, we remove the step embeddings to ensure the flexibility of steps in autoregressive modeling.
Step sampling. During training, we randomly sample the scale size in each step to enhance FlexVAR’s capability to perceive any scale. Specifically, we set the maximum number of steps to 10, fixing the scale size of the first step to 1$\times$1 and the last step to 16$\times$16 (corresponding to 256$\times$256 input images), and randomly sampling the scale sizes for the intermediate steps. Each step is dropped with a 5% probability, with a maximum of 4 steps being dropped. Thus, the number of steps during training is from 6 to 10.
During inference, we use a default of 10 steps: {1, 2, 3, 4, 5, 6, 8, 10, 13, 16} (same as VAR). Our experimental results show more steps yield better performance (Fig. 7).
@@ 4.1 Implementation details
FlexVAR tokenizer. Our scalable VQVAE tokenizer is configured with a downsampling factor of 16 and is initialized with the pre-trained weights from LlamaGen [41], the codebook size is set to 8912, and the latent space dimension is set to 32. The quantization of each scale shares the same codebook. We follow the VQVAE training recipe of LlamaGen. The training is on OpenImages [21] with a constant learning rate of $10^{-4}$, AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = 0.05, a batch size of 128, and for 20 epochs. $K$ is set to 5 by default, indicating that each latent space is randomly sampled into 5 different resolutions.
FlexVAR transformer. We provide FlexVAR in three scales, with detailed configurations for each scale provided in Tab 1. FlexVAR is trained on the ImageNet-1K 256$\times$256 using 80GB A100 GPUs. The training process employs the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay rate of 0.05. The learning rate is set to 1e-4, with the training epochs varying between 180 and 350 depending on the model scale.
@@
Table 1: Architectural design and training epochs of FlexVAR.
Table 2: Generative model comparison on class-conditional ImageNet 256$\times$256. Metrics include Fréchet inception distance (FID), inception score (IS), precision (Pre) and recall (rec). Step: the number of model runs needed to generate an image. Time: the relative inference time of VAR-$d30$ [44]. We present models with a size $\leq$ 1B.
@@
BigGAN [3] 6.95 224.5 0.89 0.38 112M 1 --
GigaGAN [20] 3.45 225.5 0.84 0.61 569M 1 --
StyleGan-XL [40] 2.30 265.1 0.78 0.53 166M 1 0.2
@@
MaskGIT [5] 6.18 182.1 0.80 0.51 227M 8 0.4
RCG (cond.) [24] 3.49 215.5 -- -- 502M 20 1.4
@@
VQGAN-re [12] 18.65 80.4 0.78 0.26 227M 256 7
RQTran. [22] 13.11 119.3 -- -- 821M 68 -
LlamaGen-XL [41] 2.62 244.08 0.80 0.57 775M 256 27
AiM [23] 2.56 257.2 0.81 0.57 763M 256 12
@@
4.2 State-of-the-art image generation
We compare FlexVAR with existing generative methods on the ImageNet-1K benchmark, including GAN, diffusion models, random-scan, raster-scan, and scaling-scan autoregressive models. As shown in Tab. 2, 3.
Overall comparison on ImageNet 256$\times$256. To ensure a fair comparison, we present models with a size smaller than 1B in Tab. 2. Our FlexVAR achieves state-of-the-art performance in all generative methods, and performs remarkably well compared to the VAR counterparts. Specifically, we achieve -0.45, -0.56, and -0.12 FID improvement compared with VAR at different model sizes.
Zero-shot inference with more steps. We use 13 steps for image generation without training, as shown in the last row of Tab. 2. FlexVAR can flexibly adopt more steps to improve image quality. By using 13 inference steps, FlexVAR further enhances the performance to 2.08 FID and 315 IS, manifesting strong flexibility and generalization capabilities. The specific steps design is detailed in the Appendix.

### v2

@@ Abstract
This work challenges the residual prediction paradigm in visual autoregressive modeling and presents FlexVAR, a new Flexible Visual AutoRegressive image generation paradigm. FlexVAR facilitates autoregressive learning with ground-truth prediction, enabling each step to independently produce plausible images. This simple, intuitive approach swiftly learns visual distributions and makes the generation process more flexible and adaptable. Trained solely on low-resolution images ($\leq$ 256px), FlexVAR can: (1) Generate images of various resolutions and aspect ratios, even exceeding the resolution of the training images. (2) Support various image-to-image tasks, including image refinement, in/out-painting, and image expansion. (3) Adapt to various autoregressive steps, allowing for faster inference with fewer steps or enhancing image quality with more steps. Our 1.0B model outperforms its VAR counterpart on the ImageNet $\times256$ benchmark. Moreover, when zero-shot transfer the image generation process with 13 steps, the performance further improves to 2.08 FID, outperforming state-of-the-art autoregressive models AiM/VAR by 0.25/0.28 FID and popular diffusion models LDM/DiT by 1.52/0.19 FID, respectively. When transferring our 1.0B model to the ImageNet $\times512$ benchmark in a zero-shot manner, FlexVAR achieves competitive results compared to the VAR 2.3B model, which is a fully supervised model trained at $\times512$ resolution.
@@
Figure 1: Generated samples from FlexVAR-$d24$ (1.0B). FlexVAR generates images with various resolutions and aspect ratios, even though it is trained with a resolution of $\leq$ $\times256$.
@@ 1 Introduction
Figure 2: Comparison between VAR var and our FlexVAR. VAR predicts the GT1 in step 1 and then predicts the residuals relative to the GT in all subsequent steps. Our FlexVAR predicts the GT at each step.
Autoregressive (AR) models aim to learn the probability distribution of the next token, offering great flexibility by generating tokens of any length. This design brings significant advancements in the field of Natural Language Processing (NLP), demonstrating satisfactory generality and transferability gpt3 ; chatgpt ; gpt4 . Concurrently, the computer vision field has been striving to develop large autoregressive models lu2022unified ; lu2024unified ; bai2024sequential ; team2024chameleon ; luo2024open ; ren2025videoworld . These models employ visual tokenizers to discretize images into a series of 1D tokens razavi2019generating ; lee2022autoregressive ; yu2021vector ; zheng2022movq ; llamagen or 2D scales var ; zhang2024var ; tang2024hart ; ren2024m ; li2024imagefolder and then utilize AR to model the next unit. However, these image autoregressive models typically output images at a single resolution, the flexibility of AR has not yet been realized.
Recently, in image generation, VAR var has pioneered scale-wise autoregressive modeling, completing image autoregression based on 2D sequences. This approach predicts the next scale rather than the next token, thereby preserving the 2D structure of images and mitigating the issue of limited receptive fields in 1D causal transformers. Specifically, VAR predicts the ground-truth (GT) [^1] To avoid confusion, we use ground-truth (GT) to represent image latent feature, and residual to represent residual latent feature. of the smallest scale in the first step. Subsequently, at each step, it predicts the residuals of the current scale and the prior one. Finally, the outputs of each scale are upsampled to a uniform size and undergo weighted summation to generate the final output, as illustrated in Fig. 2(a). Successors zhang2024var ; tang2024hart ; ren2024m ; li2024imagefolder have all adopted the residual design, assuming it to be effective. Although this technique achieves commendable performance, it encounters a primary challenge: The residual prediction relies on a rigid step design, restricting the flexibility to generate images with varying resolutions and aspect ratios, thus limiting the adaptability and flexibility of image generation. Meanwhile, residuals at different scales often lack semantic continuity, and this implicit prediction approach may limit the model’s capacity to represent diverse image variations.
@@ 2 Related Work
VQ-VAE vqvae introduces a groundbreaking two-stage image generation paradigm: (1) encoding the image into a latent space and quantizing it to the nearest code in a fixed-size codebook; (2) modeling the discretized code using PixelCNN pixelcnn , which predicts the probability distribution of each code in raster scan order. This two-stage paradigm has laid the foundation for many subsequent works.
@@
Raster-scan Manner Building on the aforementioned foundation, esser2021taming ; ramesh2021zero perform autoregressive learning in latent space with Transformer architecture. VQVAE-2 vqvae2 and RQ-Transformer lee2022autoregressive use extra scales or stacked codes for next-image-token prediction. These works further advance the field and achieve impressive results. Recently, llamagen ; liu2024lumina utilize a GPT-style next-token-prediction strategy to achieve high-quality image generation. he2024mars further improves this paradigm by introducing a mixture of autoregressive models, while aim incorporates Mamba structure gu2023mamba to accelerate image generation. xie2024show ; zhou2024transfusion ; gu2024dart combine diffusion processes into autoregressive modeling to address the information loss caused by quantization, which potentially degrades the quality of generated images.
@@
Random-scan Manner Masked-prediction models learn to predict masked tokens in a BERT-style manner devlin2018bert ; he2022masked ; bao2021beit . They introduce a bidirectional transformer that predicts masked tokens by attending to unmasked conditions, thus generating image tokens in a random-scan manner. This approach enables parallel token generation at each step, significantly improving inference efficiency. Specifically, chang2022maskgit ; chang2023muse apply masked-prediction models in class-to-image and text-to-image generation, respectively. MagViT series yu2023magvit ; luo2024open adapts this approach to videos by introducing a VQVAE for both images and videos. NOVA deng2024autoregressive first predicts temporal frames and then predicts spatial sets within each frame to achieve high-quality image/video generation.
@@
Scaling-scan Manner VAR var establishes a new generation paradigm that redefines autoregressive learning on images from next-token-prediction to next-scale-prediction. VAR in parallel predicts image tokens at one scale, significantly reducing the number of inference steps. Following VAR, VAR-CLIP zhang2024var achieves text-to-image generation by converting the class condition token into text tokens obtained from the CLIP. In terms of operational efficiency, code introduces an efficient decoding strategy, ren2024m incorporates linear attention mechanisms to accelerate image generation, and li2024imagefolder designs a lightweight image quantizer, significantly reducing training costs. Regarding generation quality, tang2024hart ; Ren2024FlowAR optimize image details by using continuous tokenizers in combination with flow matching or diffusion model. Infinity han2024infinity redefines the visual autoregressive model under a bitwise token prediction framework, remarkably enhancing generation capability and detail.
@@ 3.1 Scale-wise Autoregression (Preliminary)
Previous approaches var ; zhang2024var ; code ; ren2024m ; li2024imagefolder ; tang2024hart typically follow a residual prediction paradigm. They only regress the ground-truth at the first scale ($g_{1}$), while at subsequent $i^{th}$ scale, the residual between the preceding scale ($g_{i-1}$) and current scale ($g_{i}$) is predicted. We formulate residuals ($\{r_{i}\}^{n}_{i=2}$) as:
@@
here $\mathrm{Upsample^{i}}$ represents upsample $g_{i-1}$ to the $i^{th}$ scale. The autoregressive likelihood is:
@@
attention mechanisms (e.g., Transformer vaswani2017attention ) are utilized to instantiate this modeling. During the $i^{th}$ autoregressive step, all preceding residuals are merged in the autoregressive model, which then predicts the probability distribution of $r_{i}$. The $h_{i}\times w_{i}$ tokens in $r_{i}$ are generated in parallel, conditioned on all preceding units. Thus, image token maps can be redefined as: $T=\{g_{1},r_{2},r_{3}...,r_{n}\}$. Finally, each image token map in $T$ is upsampled to $\mathcal{R}^{h_{n}\times w_{n}}$ and summarized for image generation.
@@ 3.3 Quantize & reconstruct images at various scales
Encoding. Given an input image $I\in\mathcal{R}^{H\times W}$, an autoencoder $\mathcal{E}(\cdot)$ esser2021taming is used to convert $I$ into latent space $f$:
@@
here $h=\frac{H}{16},w=\frac{W}{16}$. We then downsample $f$ at $K$ random scales to obtain multi-scale latent features $\mathcal{F}=\{f_{1},f_{2},...,f_{K}\}$. $f_{k}$ represents represents the latent feature of the $k^{th}$ downsample from $f$. $f_{K}$ matches the original resolution of $f$.
@@
Decoding. The multi-scale images $\hat{\mathcal{I}}=\{\hat{I_{1}},\hat{I_{2}},...,\hat{I_{K}}\}$ are reconstructed using the decoder $\mathcal{D}(\cdot)$ esser2021taming given $q^{(k,i,j)}$. We follow Llamagen llamagen to adopt the same loss functions ($\mathcal{L}_{vae}$) to train $\{\mathcal{E},\mathcal{Q},\mathcal{D}\}$ at each scale without special design. Therefore, the final loss function can be formulated:
@@ 3.4 Visual autoregressive modeling without residual
In our FlexVAR, we design a 2D scalable PE ($\mathcal{P}\in\mathcal{R}^{d\times 2h\times 2w}$) adding to the h-w coordinates. It contains $2h\times 2w$ learnable queries with $d$ channels. At the ith step, $\mathcal{P}$ is upsampled/downsampled to match the scale of $g_{i}$. To ensure stability during linear interpolation across various scales, we set $\mathcal{P}$ to $2\times$ the size of the max latent space in traning. $\mathcal{P}$ is initialized using 2D sin-cosine PE vit to ensure the 2D positional correlation. Additionally, we experimentally find that in our ground-truth prediction paradigm, incorporating PE for step embeddings is unnecessary (Tab. 6). Therefore, we remove the step embeddings to ensure the flexibility of steps in autoregressive modeling.
Step sampling. During training, we randomly sample the scale size in each step to enhance FlexVAR’s capability to perceive any scale. Specifically, we set the maximum number of steps to 10, fixing the scale size of the first step to $\times1$ and the last step to $\times16$ (corresponding to $\times256$ input images), and randomly sampling the scale sizes for the intermediate steps. Each step is dropped with a 5% probability, with a maximum of 4 steps being dropped. Thus, the number of steps during training is from 6 to 10. During inference, we use a default of 10 steps: {1, 2, 3, 4, 5, 6, 8, 10, 13, 16} (same as VAR). Our experimental results show more steps yield better performance (Fig. 6).
@@ 4.1 Implementation details
FlexVAR tokenizer. Our scalable VQVAE tokenizer is configured with a downsampling factor of 16 and is initialized with the pre-trained weights from LlamaGen llamagen , the codebook size is set to 8912, and the latent space dimension is set to 32. The quantization of each scale shares the same codebook. We follow the VQVAE training recipe of LlamaGen. The training is on OpenImages openimages with a constant learning rate of $10^{-4}$, AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = 0.05, a batch size of 128, and for 20 epochs. $K$ is set to 5 by default, indicating that each latent space is randomly sampled into 5 different resolutions.
@@
Table 1: Configuration of FlexVAR.
FlexVAR transformer. We provide FlexVAR in three scales, with detailed configurations for each scale provided in Tab 1. FlexVAR is trained on the ImageNet-1K $\times256$ using 80GB A100 GPUs. The training process employs the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay rate of 0.05. The learning rate is set to 1e-4, with the training epochs varying between 180 and 350 depending on the model scale.
Table 2: Generative model comparison on class-conditional ImageNet $\times256$. Metrics include Fréchet inception distance (FID), inception score (IS), precision (Pre) and recall (rec). Step: the number of model runs needed to generate an image. Time: the relative inference time of VAR-$d30$ var . We present models with a size $\leq$ 1B.
@@
BigGAN (biggan, ) 6.95 224.5 0.89 0.38 112M 1 -
GigaGAN (gigagan, ) 3.45 225.5 0.84 0.61 569M 1 -
StyleGan-XL (stylegan-xl, ) 2.30 265.1 0.78 0.53 166M 1 0.2
@@
MaskGIT (chang2022maskgit, ) 6.18 182.1 0.80 0.51 227M 8 0.4
RCG (cond.) (li2024return, ) 3.49 215.5 - - 502M 20 1.4
@@
VQGAN-re (esser2021taming, ) 18.65 80.4 0.78 0.26 227M 256 7
RQTran. (lee2022autoregressive, ) 13.11 119.3 - - 821M 68 -
LlamaGen-XL (llamagen, ) 2.62 244.08 0.80 0.57 775M 256 27
AiM (aim, ) 2.56 257.2 0.81 0.57 763M 256 12
@@
4.2 Overall Comparison
We compare FlexVAR with existing generative methods on the ImageNet-1K benchmark, including GAN, diffusion models, random-scan, raster-scan, and scaling-scan autoregressive models. As shown in Tab. 2. To ensure a fair comparison, we only present models with a size smaller than 1B. Our FlexVAR achieves state-of-the-art performance in all generative methods, and performs remarkably well compared to the VAR counterparts. Specifically, we achieve -0.45, -0.56, and -0.12 FID improvement compared with VAR at different model sizes.
4.3 Zero-shot Comparison
Zero-shot inference with more steps. We use 13 steps for image generation without training, as shown in the last row of Tab. 2. FlexVAR can flexibly adopt more steps to improve image quality. By using 13 inference steps, FlexVAR further enhances the performance to 2.08 FID and 315 IS, manifesting strong flexibility and generalization capabilities. The specific steps design is detailed in the Supplementary Material.